I need help with this problem. Thank you
a number subtracted from 13 is the quotient of -34 and 2. Find the number.

Answer by garcia_marina(3)

Let X be the number subtracted from 13. So
13-X=(-34/2)
13-X = -17
-X = -17-13=-30
X=30